The current of 4.8 A is flowing in a conductor. The number of electrons passing through any cross section per second is: 
A. 3 x 10^16
B. 3 x 10^19
C. 7.6 x 10^12
D. 76.8 x 10^19

Grade:12th pass

1 Answers

Arun
4 years ago
As Q=I X t      .....(1)
where, I=Current, t=time and Q=Total charge flowing.
Since, Q=ne     ......(2)
where n=no.of electrons and e=charge on the electron.
Given I=4.8 A
          t=1 s
As we know e=1.6X10-19coulomb.
from eq.(1) and (2) :-
I*t=n*e
n = I * t/e = 3 * 10^19